OpenStudy (anonymous):

PLEASE HELP: MEDAL/FAN You need to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. FUNCTION: A (t) = 39,145(1 + (r/12))12t

OpenStudy (anonymous):

so... what would I do?

OpenStudy (anonymous):

@mathmale @Glizzy_Nedd @undeadknight26 @UnkleRhaukus @mathhelpnow

OpenStudy (anonymous):

SO.... What I think is that I plug in the time as 1 month? and I change the rate of .03 to .25?

OpenStudy (anonymous):

so... i would get A = 39,145(1 + (0.25/12))12 A = 50, 134 Thats what i got.... but i dont think thats right...

OpenStudy (anonymous):

is the answer choice

OpenStudy (anonymous):

no... sorry

OpenStudy (anonymous):

@undeadknight26 can you help?

undeadknight26 (undeadknight26):

plug a number in for t.

OpenStudy (anonymous):

would it be 1? cuz t stands for time..... and 1 month?

undeadknight26 (undeadknight26):

yup...then 2 then 3 etc...

OpenStudy (anonymous):

the formula is A(t) = P(1 + r/n)^nt and the problem says solve for r...

undeadknight26 (undeadknight26):

Well no idea in that case o.o

OpenStudy (anonymous):

hmmm.... O.O idk

OpenStudy (phi):

is there more info to this problem ?

OpenStudy (anonymous):

Vehicle: Chevy Volt Price: $39,145

OpenStudy (anonymous):

The rest of the info is up above

OpenStudy (anonymous):

A(t) = P(1 + r/n)^nt P is principle r is rate t is time n is number of times compounded

OpenStudy (anonymous):

and it will take 45 months for it to be compounded completely...

OpenStudy (anonymous):

i think that's about it....

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

is the rate given 3%?

OpenStudy (anonymous):

.03/12=.0025

OpenStudy (phi):

The question is still unclear. It appears they want to change the compounding to 4 times a month (which is 48 times per year) and then find the new rate? we can change to compounding 48 times per month, by replace 12 with 48 but to solve for the rate, we need more info

OpenStudy (anonymous):

that was the original rate..... but i thought you have to find the rate in this problem.... I got this: A (t) = 39,145(1 + r )^(12(-45.5)) --- 12 and solve for r?

OpenStudy (phi):

yes, if you post it.

OpenStudy (anonymous):

I can attach my problem...

OpenStudy (anonymous):

Im stuck on problem 3...

OpenStudy (phi):

to solve for the new r, you need to fill in A(t) on the left side (10,000 ?) because you are compounding 4 times per month, change the 12 to 4*12=48 the exponent was in months. You should change that to weeks (4 wks to a month) so the exponent is -45.5*4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so -182

OpenStudy (phi):

let me check

OpenStudy (anonymous):

if the amount is compounded monthly,you can check the amount more than once but rate will remain same.

OpenStudy (phi):

you mean A (t) = 39,145(1 + (r /48))^-182 if you compound 4 times per month (= 48 times per year)

OpenStudy (anonymous):

oh....

OpenStudy (anonymous):

k so im confused on how to solve

OpenStudy (phi):

The parts 2 and 3 of this question make no sense to me. But if we do what they ask, you get -45.5 months for the answer to part 2. To do part 3, we need to fill in all the variables (except r), and then solve for r

OpenStudy (phi):

To begin, if you compound 48 times per year, you change your formula to A (t) = 39,145(1 + (r/48))^48t where t is the time in years

OpenStudy (phi):

Because the question is so goofy, all we can do is take a guess as to what they want. One guess is to solve \[ 10000 = 39145\left(1 + \frac{r}{48}\right)^{48\cdot \frac{-45.5}{12}}\\ 10000 = 39145\left(1 + \frac{r}{48}\right)^{-182} \]